# Uncertainty in latent representations of variational autoencoders optimized for visual tasks

### v2

@@ Acknowledgments
Gamma-Laplace 1 1800 0.46/$\alpha$ 1 1 2 3e-5 1e-5 10000 128
($\alpha\approx 5.315$) $\approx 0.09$
Table S1: Hyperparameter settings for the Van Hateren models
Model PC $D$ h( ) g( ) $\sigma_{obs}$ $\beta^{\prime}_{1}$ $\beta^{\prime}_{2}$ lr Epochs Batch Size
MNIST 1 5 sigmoid exp 4 1 1e-3 500 128
ChestMNIST 1 5 sigmoid exp 1 1 1e-3 500 128
cMNIST10D 1 10 exp 0.25 0.125 0.125 1e-4 500 128
cMNIST3D 1 3 exp 0.25 0.125 0.125 1e-4 500 128
cFashionMNIST 1 10 exp 0.25 0.125 0.125 1e-4 500 128
cChestMNIST 1 4 exp 0.1 0.02 0.02 1e-4 500 128
Table S2: Hyperparameter settings for the MNIST domain models
@@
Fig. S4: Morphing experiments in MNIST. a, Centers for the distribution corresponding to each digit are first found. b, Posterior widths for the morphed digits with standard VAE (left) and EA-VAE (right). Position where uncertainty is maximal is computed via quadratic fit. c, For a given window size around the middle point between the centers, the fraction of interpolations where the maximal uncertainty lies within this window is computed via b fittings.
@@
Fig. S5: Out of distribution experiments. Same as in Fig. 2d, now when training on a, cMNIST 10D, b, cFashionMNIST 10D, c, cChestMNIST, d, ChestMNIST and e, Gamma-Laplace Van Hateren ($z$-scored).
@@
Fig. S6: Behaviour of trained MLP classifier when gradually morphing between image samples from digit ’A’ to digit ’B’. Entropy of predicted probability distribution as a function of combination weight of labels for all 45 possible combinations of labels. Dashed line corresponds to the upper limit determined by the discrete uniform distribution.

### v3

@@ 3 Supplementary Tables and Figures
[…]
Table S2: Hyperparameter settings for the NIST and NIH-ChestXray14 domain models
Model PC $D$ h( ) g( ) $\sigma_{obs}$ $\beta^{\prime}_{1}$ $\beta^{\prime}_{2}$ lr wd Epochs Batch Size
MNIST 1 5 sigmoid exp 4 1 1e-3 1e-5 500 128
ChestMNIST 1 5 sigmoid exp 1 1 1e-3 1e-5 500 128
NIH-ChestXray14 1 20 sigmoid softplus 50 50 5e-4 1e-5 500 128
cMNIST10D 1 10 exp 0.25 0.125 0.125 1e-4 1e-5 500 128
cMNIST3D 1 3 exp 0.25 0.125 0.125 1e-4 1e-5 500 128
cFashionMNIST 1 10 exp 0.25 0.125 0.125 1e-4 1e-5 500 128
cChestMNIST 1 4 exp 0.1 0.02 0.02 1e-4 1e-5 500 128
MNIST-LinearDecoder 1 5 exp 0.1 0.1 5e-4 1e-5 2500 128
Fig. S1: a, Trained VAE (left) and EA-VAE (right) models for inference on natural images. The encoder’s input of size $M=1600$ is followed by two fully connected hidden layers with softplus activation function. Two separate fully connected layers of $D=1800$ units encode the mean and variance of $q_{\phi}(\mathbf{z}|\mathbf{x})$. For the VAE, one fully connected layer decodes the latent sample into a reconstructed patch. For the EA-VAE model, the described VAE was modified such that the last component of the sample of the posterior distribution passes through function g() defining the effective $s$ sample from $q_{\psi}(s|\mathbf{x})$. The posterior sampled $s$ value is globally multiplied with the decoder’s output, resulting in the final reconstructed patch. b, Trained VAE (left) and EA-VAE (right) models for inference on natural images when scaling variable assumed from a Gamma distribution. In this case, the scaling variable is inferred from a separate encoder. $k=2$ was chosen as the smallest integer number that qualitatively achieved a similar distribution to the dataset’s $\sigma_{pix}$ distribution. Computationally, the choice of a fixed integer $k$, simplifies the sampling procedure, since one can always sample from the same standard Gamma function and then reparameterize by the scale parameter. This enables simple gradient computations across the sampling process. c, Trained VAE (left) and EA-VAE (right) models for inference on MNIST classic and contrast-augmented domains (digits, chest x-rays and fashion). For the VAE, the encoder’s input of size 32x32 is followed by three convolutional hidden layers with relu activation function. Two separate fully connected layers of $D$ units encode the mean and variance of $q_{\phi}(\mathbf{z}|\mathbf{x})$. A fully connected and 3 deconvolutional layers decode the latent sample. Output from last layer passes trough h() function giving place to the reconstructed image. For the EA-VAE, the described VAE model in was modified such that separate fully connected layers encode mean and variance of a Normal distribution. Its sample passes through function g() defining the effective $s$ sample from $q_{\chi}(s|\mathbf{x})$. The posterior sampled $s$ value is globally multiplied with the decoder’s output before going through function h(), resulting in the final reconstructed image. d, Trained VAE (left) and EA-VAE (right) models for inference on NIH-ChestXray14 dataset. For the VAE, the encoder’s input of size 224x224 is followed by four convolutional hidden layers with relu activation function. Two separate fully connected layers of $D$ units encode the mean and variance of $q_{\phi}(\mathbf{z}|\mathbf{x})$. A fully connected and 4 deconvolutional layers decode the latent sample. Output from last layer passes trough h() equal to sigmoid() function giving place to the reconstructed image. For the EA-VAE, the described VAE model in was modified such that separate fully connected layers encode mean and variance of a Normal distribution. Its sample passes through function g() equal to softplus() defining the effective $s$ sample from $q_{\chi}(s|\mathbf{x})$. The posterior sampled $s$ value is globally multiplied with the decoder’s output before going through function h(), resulting in the final reconstructed image.
@@
Fig. S4: Out of distribution experiment when training a Standard VAE and EA-VAE on MNIST with a linear generative model. Posterior width and the magnitude of the scaling variable for in-distribution (ID) and out-of-distribution (OOD) examples both for standard VAE (blue) and EA-VAE (red). Top: Examples of images from the same domain (light green frames) and different domains (mustard frames) as inputs, and reconstructions by the standard VAE (second row) and EA-VAE (third row) respectively. Bottom: uncertainty quantified by the posterior width and mean of scaling variable posterior for individual within-distribution or out-of-distribution images. Architecture of Figure˜S1c was used, without the relu activations in the decoder. Hyperparameters in Table Table˜S2.
@@
Fig. S5: Out of distribution experiments. Same as in Fig. 3a, now when training on a, cMNIST 10D, b, cFashionMNIST 10D, c, cChestMNIST, d, ChestMNIST and e, Gamma-Laplace Van Hateren ($z$-scored).
Fig. S6: Morphing experiments in MNIST. a, Centers for the distribution corresponding to each digit are first found. b, Posterior widths for the morphed digits with standard VAE (left) and EA-VAE (right). Position where uncertainty is maximal is computed via quadratic fit (shown in gray curve). c, For a given window size around the middle point between the centers, the fraction of interpolations where the maximal uncertainty lies within this window is computed via b fittings.
Fig. S7: Posterior width and the magnitude of the scaling variable of standard VAE and EA-VAE models trained on ChestMNIST, for images increasingly corrupted either by Gaussian blurring (top) or additive pixel noise (bottom).
Fig. S8: Latent posterior width (mustard dots, with color matching that of the frame of average images) for uninformative images (top, mustard framed images) in the MNIST (left) and ChestMNIST (right) data sets. The distribution of noise posterior widths (std.) for in-distribution test images for standard VAE and EA-VAE are presented as violin plots. Prior uncertainty is shown as a reference (dashed line).
@@
Fig. S9: Behaviour of trained MLP classifier when gradually morphing between image samples from digit ’A’ to digit ’B’ when trained with standard VAE (blue) and EA-VAE (red) latent representations. Entropy of predicted probability distribution as a function of combination weight of labels for all 45 possible combinations of labels. Dashed line corresponds to the upper limit determined by the discrete uniform distribution.
Fig. S10: a, VAE and EA-VAE trained on Imagenet from Inception feature space. Top: The pretrained Inception network extracts for a given image, a feature vector (blue box) from which the image is then classified (lilac box). Bottom: VAE/EA-VAE is trained using as input the feature vector extracted from pre-trained fixed Inception block. b, Example Imagenet images used to train and test the VAE/EA-VAE. Left: Human-class (green label) Ball Player images were used to train. Right: Human-class Groom and Scuba Diver, as well as Object-class (orange label) Desk, Wall Clock and Bookcase were used to test out of distribution examples. c, Posterior width for in-distribution (Ball Player) and out-of-distribution (Groom, Scuba Diver, Desk, Wall Clock and Bookcase) examples when trained on Inception feature vector for Ball Player images both for standard VAE (blue) and EA-VAE (red). As a reference, a dashed black line indicates the prior uncertainty. d, VAE vs EA-VAE capability to differentiate out-of-distribution examples. Comparison of posterior width for in-distribution and out-of-distribution for a single model is measured as the $\log_{10}(\text{p-value})$ for the t-test between the distributions. Each point corresponds to one of the out-of-distribution classes, and error bars arise from the $N=10$ trained models. As a reference, a dashed black line indicates the identity.
Fig. S11: Uncertainty estimates of the Importance Weighted Autoencoder on the Van Hateren dataset. The signal means, signal variances and noise (posterior) variances for the test images are shown as in the main text in three IWAE variants, with $k=5,10,50$ samples respectively. The updated architectures do not show any different representation than the standard VAE presented in the paper.
Fig. S12: Characterization of posterior width for systematic image manipulations that affect inference uncertainty for VAE with active dropout during training in the encoder of the VAE, in comparison to standard VAE and EA-VAE. a, Latent posterior width (mustard dots, with color matching that of the frame of average image) for uninformative images (top, mustard framed images) in the MNIST dataset. The distribution of noise posterior widths (std.) for in-distribution test images for each model are presented as violin plots. Prior uncertainty is shown as a reference (dashed line). b, Posterior widths for gradual morphing between representative image samples from digit ’i’ and digit ’j’ with standard VAE (left), VAE with dropout (center) and EA-VAE (right). c, Posterior width for in-distribution (ID) and out-of-distribution (OOD) examples when trained on MNIST for standard VAE (blue), VAE with dropout (green) and EA-VAE (red). Top: Examples of images from the same domain (gray frames) and different domains (mustard frames) as inputs, and reconstructions by the standard VAE (second row), VAE with dropout (third row) and EA-VAE (fourth row) respectively. Bottom: uncertainty quantified by the posterior width for individual within-distribution or out-of-distribution images. As a reference, a dashed black line indicates the prior uncertainty.